OpenStudy (anonymous):

There are 23 coins in a bank. All the coins are dimes and quarters. The total value of the coins is $3.80. How many dimes are there? How many quarters?

OpenStudy (anonymous):

these question are looking mighty familiar

OpenStudy (anonymous):

No. of dimes (d): 10d + 25(23 - d) = 380 2d + 5(23 - d) = 76 2d + 115 - 5d = 76 3d = 39 d = 13 No. of quarters: = 23 - 13 = 10

OpenStudy (anonymous):

try \[10x+25(23-x)=380\]

OpenStudy (anonymous):

what randicandi said

OpenStudy (anonymous):

ty:D